### v1

# Where Are We? Evaluating LLM Performance on African Languages

Abstract
Africa’s rich linguistic heritage remains underrepresented in NLP, largely due to historical policies that favor foreign languages and create significant data inequities. In this paper, we integrate theoretical insights on Africa’s language landscape with an empirical evaluation using Sahara— a comprehensive benchmark curated from large-scale, publicly accessible datasets capturing the continent’s linguistic diversity. By systematically assessing the performance of leading large language models (LLMs) on Sahara, we demonstrate how policy-induced data variations directly impact model effectiveness across African languages. Our findings reveal that while a few languages perform reasonably well, many Indigenous languages remain marginalized due to sparse data. Leveraging these insights, we offer actionable recommendations for policy reforms and inclusive data practices. Overall, our work underscores the urgent need for a dual approach—combining theoretical understanding with empirical evaluation—to foster linguistic diversity in AI for African communities.
1 Introduction
The digital landscape of Africa is as diverse as its cultures and languages, yet the continent’s rich linguistic heritage remains largely underrepresented in NLP research. Despite significant transformations over the past decade driven by advances in large language models (LLMs) and a growing interest in linguistic diversity, most efforts have concentrated on a handful of widely spoken languages such as Swahili, Afrikaans, and Hausa, while many Indigenous languages continue to be sidelined due to the paucity of accessible, high-quality datasets. This oversight, deeply rooted in historical language policies that have neglected robust data collection and accessibility, has far-reaching implications for the development of language technology across Africa Eberhard et al. (2021). Recent initiatives, including the work of Masakhane Dione et al. (2023b); Adelani et al. (2021c, 2023, 2024b); Bandarkar et al. (2024a) and the efforts of platforms like HuggingFace, have begun to address these challenges by curating language-specific datasets and adapting transfer learning techniques for low-resource African languages. Models such as Cheetah Adebara et al. (2024a), Serengeti Adebara et al. (2023), Toucan Elmadany et al. (2024), Afri-XLMR Alabi et al. (2022), AfriTeVa Oladipo et al. (2023); Jude Ogundepo et al. (2022), mBERT Devlin et al. (2018), XLM-R Conneau et al. (2020), and LLaMA have leveraged multilingual data to enhance performance on these languages, yet progress remains uneven—highlighting that the advancements achieved are primarily concentrated on a select few.
In this study, we undertake a comprehensive empirical evaluation of leading large language models (LLMs) on an extensive benchmark that we collect using mostly existing datasets to unravel how current African language policies, by dictating data availability and accessibility, directly shape model performance across the continent’s diverse linguistic spectrum. By systematically analyzing the distribution of existing datasets across languages and performance patterns among various LLMs, we not only pinpoint which languages benefit from ample data resources but also illuminate the underlying factors that contribute to their relative success. Our investigation reveals that the disparities in NLP outcomes are closely tied to policy-driven data inequities, offering concrete evidence of the need for more inclusive, forward-thinking language policies. These insights provide actionable recommendations for enhancing dataset creation and model training, ultimately aiming to bridge the digital divide and foster linguistic diversity in artificial intelligence for African communities.
Overall, we make the following contributions: (1) Benchmark with wide, diverse coverage. We introduce Sahara, a comprehensive benchmark assembled from large-scale, publicly accessible, and inclusive datasets that capture the rich linguistic diversity of Africa. This benchmark enables a systematic evaluation of data availability and model performance across a broad range of African languages. (2) Dynamic leaderboard. We develop a dynamic leaderboard built on best design practices, providing a transparent and continually updated platform to track progress and benchmark innovations in African NLP. (3) Empirical evaluation. We perform an extensive empirical assessment of a wide spectrum of models thereby establishing robust baselines and revealing performance disparities linked to data resource availability. (4) Policy-driven and actionable insights. We uncover clear links between current language policies, data availability, and model performance, demonstrating how policy-induced data inequities shape the effectiveness of language models across Africa. Our analysis provides recommendations for future research and development by informing policy reforms that foster more inclusive language technologies tailored to the needs of African communities. We now discuss language polices in Africa.
2 Language Policies in Africa
(a) Language policies for education across Africa.
(b) Official national language policies across Africa.
(c) Official regional language policies across Africa. None means no additional policy is available on a regional level.
(d) We collect  Saharato empirically evaluate LLM performance on African languages, allowing us to demonstrate with evidence how current language policies directly impact progress in the field.
Figure 1: Maps of Africa showing the languages covered in this work and the language policies across Africa. The term Both in maps (a), (b), and (c) refer to Indigenous and foreign languages combined.
Across Africa, the predominant response to the continent’s multilingual landscape has been the adoption of a foreign language for official functions. In some instances, select Indigenous African languages receive official recognition at regional or national levels or within educational contexts Petzell (2012); Foster (2021); Ouane and Glanz (2010). Figures 1(a), 1(b), and 1(c) illustrate the distribution of language policies across the continent. [^1] We designed the maps in these figures based on information derived from Ethnologue. We checked the national, regional, institutional, and educational languages for each country. For example, in Nigeria, English is the official language, with only three out of 512 Indigenous languages recognized at the regional level. Similarly, Ghana designates English as its official language while also recognizing ten of its 73 Indigenous languages for institutional use. In Tanzania, Swahili is the sole Indigenous official language among 118 languages, alongside English. Kenya grants official status to $12$ of its $61$ languages, while South Africa recognizes $12$ out of $20$ Indigenous languages as institutional languages Adebara and Abdul-Mageed (2022).
Even when Indigenous languages are granted official status alongside foreign languages, they often serve a symbolic rather than functional role. For instance, although Kiswahili is recognized as an official language by the African Union, its website and official document releases remain in English and French. A similar pattern emerges in education systems: where Indigenous languages are used, their role is typically confined to early childhood education and is usually paired with a foreign language rather than serving as the sole medium of instruction Petzell (2012); Foster (2021); Ouane and Glanz (2010). These language policies significantly shape language use across various domains, including newspapers, radio, television, and social media. Due to the dominance of foreign languages in official contexts, major newspapers and government publications are predominantly produced in languages such as English, French, and Portuguese, thereby limiting access for speakers of Indigenous languages ResCue and Agbozo (2021). Although some Indigenous languages are used on television and radio, foreign languages tend to dominate national broadcasts, especially in political discourse and formal news reporting Cheo et al. (2023); Myers and Harford (2020); National Communications Authority (2016).
Social media presents a more flexible environment where users can communicate in Indigenous languages; however, platform support remains uneven Sunday et al. (2018); Molale and Mpofu (2023); ResCue and Agbozo (2021). Many African languages are underrepresented in digital spaces, lacking features such as text input options, spell checkers, or automated translation services. Consequently, users often resort to code-switching between Indigenous and dominant foreign languages, reinforcing existing linguistic divides. These trends underscore how official language policies shape broader language use, ultimately contributing to the marginalization of Indigenous languages in both public and digital communication spheres.
The analysis of language policies in Africa reveals how historical choices have led to significant data imbalances for Indigenous languages. Motivated by this, we built our benchmark, Sahara to evaluate how these policies directly impact data availability, coverage, and the performance of LLMs on African languages. In the following section, we detail the construction of  Sahara and demonstrate its role in linking language policy to empirical outcomes in NLP.
3 Related Work
The development of NLP for African languages is hindered by data scarcity, policy-induced disparities, and suboptimal model performance. Recent initiatives have sought to mitigate these challenges by curating diverse datasets, rigorously evaluating multilingual models, and advocating for more inclusive language policies. For example, several initiatives have contributed to expanding datasets for African Yimam et al. (2020); Muhammad et al. (2022); Aliyu et al. (2022); Muhammad et al. (2023b, a); Ilevbare et al. (2024); Elmadany et al. (2024). Some notable ones like Adelani et al. (2021c, 2022b) introduced MasakhaNER, a named entity recognition dataset for 10 African languages, highlighting the limitations in existing resources. Similarly, IROKOBench Adelani et al. (2024b) evaluated large language models (LLMs) on African languages, revealing stark performance gaps between high- and low-resource languages. Adebara et al. (2022) developed AfroLID, a neural language identification model covering 517 African languages, showing that most languages lack NLP datasets beyond language identification.
Cluster Task Identifier #AL #DS
Classification Cross-Lingual Natural Language Inference Adelani et al. (2024b) xnli $16$ $1$
Language Identification Adebara et al. (2022) langid $517$ $1$
News Classification Azime and Mohammed (2021); Niyongabo et al. (2020); David (2020a) news $4$ $3$
Sentiment Analysis Diallo et al. (2021); Oyewusi et al. (2020); Shode et al. (2022a) sentiment $3$ $3$
Topic Classification Hedderich et al. (2020) topic $2$ $1$
Generation Machine Translation Adelani et al. (2022a); Reid et al. (2021a); Ogueji and Ahia (2019); Akera et al. (2022) mt $29$ $5$
Paraphrase Scherrer (2020) paraphrase $4$ $1$
Summarization Hasan et al. (2021); Adebara et al. (2024b) summary $10$ $2$
Title Generation Hasan et al. (2021); Adebara et al. (2024b) title-gen $10$ $2$
MCCR General Knowledge Adelani et al. (2024b) mmlu $16$ $1$
Mathematical Word Problems Adelani et al. (2024b) mgsm $16$ $1$
Reading Comprehension Bandarkar et al. (2024a) reading-comp $25$ $1$
Context-based Question Answering Clark et al. (2020) squad-qa $1$ $1$
Tokens NER Adelani et al. (2021a, 2022c); Eiselen (2016); Alabi et al. (2020a); Pan et al. (2017) ner $27$ $5$
Phrase Chunking Eiselen (2016) chunking $8$ $1$
POS Tagging Onyenwe et al. (2018a, 2019) pos $1$ $1$
Total $\bf 16$ $\bf 517$ $\bf 30$
Table 1: Descriptive statistics of languages and dataset diversity in Sahara across different tasks and task-clusters. #AL: African Languages, #DS: datasets, MCCR: multiple-choice, comprehensive and reasoning task-cluster.
Language policies play a crucial role in determining resource availability. Adebara and Abdul-Mageed (2022); Hershcovich et al. (2022) emphasized the sociolinguistic factors affecting NLP development, noting that languages with official status receive more institutional support, fostering greater digital representation. Bird (2020) further critiqued the Western-centric approach in speech and language technologies, calling for AI models that align with Indigenous linguistic contexts. Model performance in African languages remains uneven, largely due to the dominance of high-resource languages in multilingual training. Conneau et al. (2018) introduced XLM-R, demonstrating the limitations of cross-lingual transfer learning for underrepresented languages. Meta AI’s No Language Left Behind (NLLB) project Costa-jussà et al. (2022) attempted to improve multilingual models for low-resource languages, but performance remained inconsistent, particularly in generation tasks. Despite these advances, African NLP still requires increased investment in dataset curation, inclusive policy frameworks, and model development tailored to the linguistic diversity of the continent. We now introduce Sahara.
4 [Uncaptioned image] Sahara Benchmark
Our objective is to create a comprehensive benchmark for African NLP that enables (i) analysis of existing language resources and (ii) the assessment of language models and tracking of the progress of African NLP. To achieve this goal, we develop Sahara, adhering to several key design principles that we will now elucidate. Sahara establishes a comprehensive and adaptable benchmark for African NLP using publicly available datasets. Figure 1(d) shows coverage of Sahara.
Wide and Diverse Coverage.
Sahara is the most comprehensive benchmark for African NLP, covering a vast range of languages. We achieve this by collecting high-quality, publicly available datasets from as many African languages as possible, ensuring easy accessibility for researchers evaluating their models. As a result, Sahara supports $517$ languages across various tasks, making it the most extensive and representative benchmark for African NLP. It encompasses languages from $50$ out of $54$ African countries and includes data written in five different scripts: Arabic, Coptic, Ethiopic, Latin, and Vai, spanning five language families across the continent.
Tasks and Task Clusters.
Sahara is designed to support a broad spectrum of NLP tasks, which we systematically organize into coherent task clusters. The inclusion of well-defined, challenging clusters enables a more meaningful assessment of a model’s capabilities and allows researchers to evaluate performance on specific clusters. As shown in Table1, Sahara comprises four clusters: text classification, text generation, multiple-choice, comprehensive and reasoning (MCCR), and token-level classification. For further details on the datasets used for each task, please refer to Appendix A.
Cluster Identifier Metric SLMs LLMs Closed LLMs
S1 S2 S3 S4 L1 L2 L3 L4 L5 L6 L7 CL1 CL2
Classification xnli Acc. 32.04 44.44 40.56 47.78 54.26 46.48 61.67 68.52 65.00 47.04 42.96 74.41 82.65
langid Acc. 0.24 0.59 0.70 1.89 1.11 2.34 2.21 2.45 8.69 17.64 15.45 20.56 13.34
news Acc. 13.76 12.96 2.00 2.77 42.86 2.66 18.52 54.16 27.27 75.83 7.85 58.33 67.85
sentiment Acc. 45.65 22.25 30.30 14.58 45.00 7.14 33.92 25.33 49.25 69.69 25.33 56.33 56.45
topic Acc. 36.10 34.28 17.27 15.80 57.77 10.70 38.09 42.63 78.43 67.43 15.38 57.70 79.99
Avg. 25.56 22.90 18.17 15.36 40.20 13.86 30.88 38.62 45.73 55.53 21.39 53.47 60.06
Generation mt Bleu 1.82 7.31 0.04 3.66 6.18 5.26 6.91 5.03 17.48 13.80 5.57 9.99 18.34
paraphrase Bleu 13.17 16.27 10.05 7.12 19.14 8.05 6.14 9.83 24.67 37.37 8.22 17.25 37.00
summary Rouge 7.55 9.98 2.00 9.13 11.18 9.56 2.00 14.69 13.24 11.01 14.45 9.10 12.35
title-gen Bleu 0.32 1.95 0.84 2.09 4.01 0.26 1.68 3.95 14.68 4.81 2.86 4.28 13.55
Avg. 5.72 8.88 3.23 5.50 10.13 5.78 4.18 8.38 17.52 16.75 7.78 10.16 20.31
MCCR mmlu Acc. 27.59 27.59 36.30 38.70 61.85 34.26 69.26 76.67 71.11 61.85 41.48 70.00 75.00
mgsm ExctM 1.94 1.94 1.48 5.19 8.24 1.11 12.04 15.56 17.04 10.00 5.19 16.98 22.34
rdng-comp Acc. 24.65 33.47 30.83 30.90 40.28 33.46 47.01 49.23 51.03 40.40 38.33 78.54 85.19
squad-qa F1 59.36 69.46 42.80 46.90 73.06 29.68 43.67 79.43 87.47 77.35 47.58 58.67 55.99
Avg. 28.39 33.12 27.85 30.42 45.86 24.63 43.00 55.22 56.66 45.90 33.15 56.05 59.63
Tokens ner Acc. 17.77 27.90 19.65 16.06 29.37 0.56 22.23 49.71 37.91 42.82 7.20 46.70 16.94
chunking Acc. 5.02 11.02 6.92 4.90 21.61 0.05 12.57 18.15 15.15 19.42 2.82 26.60 18.21
pos Acc. 6.34 6.96 2.80 5.12 11.13 1.31 9.05 15.09 0.06 10.10 3.46 11.02 11.04
Avg. 9.71 15.29 9.79 8.69 20.70 0.64 4.62 27.65 17.71 24.11 4.49 28.11 15.40
Overall Avg. 17.34 20.05 14.76 14.99 29.22 11.23 23.17 32.47 34.40 35.57 16.70 36.94 38.85
Table 2: Few-shot evaluation across task datasets. The highest score for each individual task is in bold green, while the best average score across all tasks is in bold orange. S1Llama3.2 (1B) S2 Llama3.2 (3B) S3 Gemma2 (2B) S4 Phi3.5 (3.8B) L1LLama3.1 (8B) L2 Aya23 (8B) L3 Gemma2 (9B) L4 Gemma2 (27B) L5 Llama3.1 (70B) L6 CommadR+ (104B) L7 Aya23 (35B) CL1 GPT4o CL1 Claude-3.5
Modular Public Leaderboard.
We implement a user-friendly leaderboard for model evaluation on Sahara, hosted on Hugging Face (HF) Spaces. Evaluated models must be publicly available in HF repositories to ensure seamless integration. When submitting a model, users are required to provide detailed metadata—including whether the model is pretrained, further pretrained, or instructed—and then submit an evaluation request. The submission enters a processing queue, and once evaluation is complete, the results are automatically displayed on the leaderboard. Our team will actively maintain the leaderboard by continuously integrating new publicly available tasks and datasets to keep it up-to-date. Sahara thus serves as a valuable tool for assessing model performance in African NLP, accelerating the development of high-performance models, and contributing to the field’s growth. In addition to evaluating models on the entire benchmark, users have the flexibility to assess performance on specific task clusters, encouraging the development of specialized models tailored to address the unique challenges of African NLP.
5 Experiments and Results
5.1 Experimental Setup
To provide an overview of the state of African NLP, we evaluated the performance of popular instruction-based LLMs in a few-shot setting using the Sahara benchmark. The models tested include Phi-3.5 Abdin et al. (2024), Gemma-2 Team et al. (2024), Llama-3.1 Dubey et al. (2024), Llama-3.2, Aya-23 Aryabumi et al. (2024), and command-r-plus from open LLMs, as well as GPT-4o and Claude-3.5-Sonnet from closed LLMs. We also considered variations of these models based on their size, categorizing the LM as small, SLM for those with fewer than 8B parameters and large, LLM for those with more than 8B parameters to analyze the effect of model size on performance. For clearer visualization, we aggregated the test datasets for each task cluster identified in Appendix A. We randomly chose a sample of $600$ data points for each task for a few-shot evaluation. The evaluation used well-crafted prompts that are both concise and straightforward, with examples of these prompts provided in Appendix C1. For tasks available in the lm-eval tool Gao et al. (2024), we use the tool for evaluating the open-source LMs.
5.2 Results and Discussion
We evaluate 13 LMs across four task clusters, maintaining a consistent assessment approach while varying evaluation metrics. Depending on the task, we employ F1, Accuracy, Bleu, and RougeL scores.
As shown in Table 2, in the classification tasks, the closed models Claude-3.5-Sonnet and GPT-4o consistently led in performance, followed by larger LLMs such as Llama3.1-70B and CommandR-Plus-104B. In the language identification task, most models struggled to grasp the task effectively, with only the larger LLMs demonstrating some understanding, suggesting that model scale is crucial for achieving higher accuracy across various multilingual tasks. In the generation task cluster, the larger models performed as expected, with Claude-3.5 and CommandR-Plus leading. There was a noticeable performance increase across all models in the paraphrase task, while the summarization task exhibited a more balanced performance among all models. Considering their size, Llama3.2-3B and Llama3.1-8B performed well within this task cluster. In the MCCR tasks, Llama3.1-70B performed competitively against the closed Claude-3.5 model. In the MGSM task, which involves solving mathematical word problems, most open-source models performed poorly, while Claude-3.5 excelled, achieving an exact match score of 22.34%. In token-level tasks, GPT-4o and CommandR+ consistently delivered strong performances, whereas the Aya-23 model family struggled to grasp the tasks.
The evaluation indicates a landscape of competitive and specialized performance among multilingual and African LLMs, with certain models standing out in some of the tasks. In few-shot evaluation, the closed models followed by Llama3.1-70B and Command-R-Plus, performed significantly better than the other models across multiple tasks. Given their size, the Llama models, especially Llama3.1-8B, perform better than the other models. Additionally, translating into English proved easier than translating from English, reflecting the impact of language representation and data availability on model efficacy (especially at the target side).
Figure 2: Distribution of languages across different clusters. Each bar represents the number of datasets a language appears in, categorized into the clusters, with some languages appearing in all four clusters while others are only present in one or two.
Which LM should I use for reasoning tasks?
In a limited resource setting, for mathematical QA tasks that fall in the MCCR cluster, Phi3.5 Abdin et al. (2024) ranks top amongst 4 other SLMs for 16 African Languages. In the LLMs category, Gemma 27B has better performance on a higher number of African languages. For reading comprehension tasks, Llama 3.2-3B Dubey et al. (2024) is most suitable in constrained compute resource settings while in scenarios without limits in compute resources, Llama 3.1-70B gives better performance.
Which African language do LMs perform best on?
Most of lama3 models perform best on Swahili. It ranks first in two models and second in one of three models tested. In the Gemma2 models, Aya-x and CommandR+, Swahili and Afrikaans are generally well known. We believe this is due to the availability of data for both languages, with both languages being among the highest resourced languages in Africa.
6 Data-Performance-Policy Analysis
To understand the relationship between policy, data availability, and performance, we carry out a number of analyses inspired by Sahara  data and empirical results on it.
(a) Generation tasks.
(b) Classification tasks.
(c) MT tasks.
(d) MT tasks grouped.
Figure 3: Distribution of best model performance (Claude 3.5-Sonnet) on different task clusters across African languages in downstream data.
Data Availability and Diversity.
As shown in Figure 2, among the $517$ languages in our benchmark, only $45$ have more than one dataset, while the majority are represented solely by language identification data. Amharic leads with $11$ unique datasets across all clusters, followed by Yorùbá with ten and Hausa with nine. A total of six languages have only one dataset beyond language identification, while the remaining $36$ have between two and eight datasets each. It is important to note that the disparity in dataset availability is closely linked to language policy. The $45$ languages with multiple datasets generally hold official status or are spoken by a significant majority in their respective countries. Government policies that designate certain languages as official often yield greater institutional support, driving increased documentation, education, and media presence. For example, Amharic’s official status in Ethiopia contributes to its relatively rich dataset availability, while Yorùbá and Hausa, both widely used in Nigeria for regional governance and media, similarly benefit from enhanced linguistic resources.
Conversely, languages lacking official recognition or widespread institutional support receive minimal investment in digital and linguistic infrastructure. The dominance of a few languages within policy frameworks perpetuates data scarcity for many Indigenous African languages, thereby limiting their inclusion in NLP advancements and reinforcing digital inequalities. Addressing this imbalance requires policy shifts that prioritize multilingual data collection, increased funding for low-resource language research, and the integration of diverse languages into education, media, and digital communication.
Data Quality.
Most of the downstream task datasets are limited to simple tasks that require merely an atomic class label at the word, sentence, or document level (e.g., sentiment analysis, named entity recognition, topic classification). Furthermore, many of these datasets are translations of existing resources or content from high-resource languages like English and French, and they do not always reflect authentic language usage within the respective communities. For instance, AfriXLNI, AfriMMLU, and AfriMGSM Adelani et al. (2024b) were created by translating XLNI, MMLU, and MGSM into African languages. In these datasets, the frequent use of borrowed words and phonologized variants of foreign terms Adebara and Abdul-Mageed (2022) highlights the challenges posed by cross-lingual concept gaps. When translations are applied to classification tasks, misalignments in the attached labels can occur because labels are not always equivalent across languages. Some labels may not exist in the target language or may have restricted usage. For example, while English possesses a productive set of adjectives, languages like Yorùbá have a more limited adjective inventory that is context-dependent, meaning that an English adjective may be translated as a noun or verb in Yorùbá rather than as an adjective.
Model Performance
Here, we analyze performance of our best model, Claude 3.5-Sonnet. As shown in Figure 3, the model faces varying degrees of difficulty when handling African languages. Sub-figure 3(a) demonstrates that for generation tasks, even the strongest models struggle to produce high-quality text in African languages, with most scores falling below 15 BLEU/ROUGE. In contrast, sub-figure 3(b) shows that while many languages achieve relatively strong classification performance (over 80% accuracy), a substantial portion still falls into the 60%–80% range, underscoring a clear performance gap. Finally, sub-figures 3(c) and 3(d) reveal that translation into English is notably easier than translation out of English or French into African languages, where BLEU scores drop from an average of around $19$ down to nine or $11$, respectively. This finding suggests that even the best model struggles more with producing fluent, faithful output in African languages than with understanding them—underscoring the urgent need for better data and model development tailored to these languages. Refer to Appendix E for a similar analysis leveraging our empirical language identification evaluation.
Apart from our empirical analyses, the limited availability and diversity of datasets for African languages in NLP benchmarks has significant implications for developing robust language technologies. Limited Model Performance and Generalization – Models trained on a single dataset per language risk overfitting to specific domains or styles, reducing their ability to generalize across different contexts, which is especially problematic for African languages with high dialectal variation. Bias Toward Well-Resourced Languages – Languages with more datasets naturally perform better in multilingual models, reinforcing disparities and leading to poor performance for underrepresented African languages. Restricted Downstream Applications – Many NLP applications, such as machine translation and question answering, require diverse datasets; if a language is represented only by language identification data, it cannot support the development of more complex language technologies. Challenges in Transfer Learning – Transfer learning techniques, like multilingual fine-tuning, rely on cross-lingual similarities and sufficient data; when African languages suffer from sparse training data, they are unable to effectively leverage knowledge transfer from resource-rich languages.
7 Recommendations
Policy determines how languages are used across various domains, influencing their visibility and viability in digital spaces. Currently, the limited online presence of many Indigenous African languages restricts the availability of training data, which is crucial for building effective NLP models. We provide the following recommendations:
Policy matters. Governments and policymakers should enact national language policies that mandate the inclusion of Indigenous and underrepresented languages across digital infrastructure, education, and media. Countries with multilingual populations should adopt policies that support the systematic documentation, transcription, and digitization of African languages to enhance their presence in AI applications. Such efforts will enrich the diversity of available data, paving the way for the development of more robust and inclusive models.
Expand data collection and annotation. Given that only $45$ of the $517$ languages in the benchmark have more than one dataset based on our collection, it is essential to increase the availability of high-quality data for underrepresented African languages across a broad range of downstream tasks. Moreover, prioritizing community-driven data annotation efforts will be crucial in ensuring both linguistic diversity and accuracy, especially for languages that lack institutional support.
Develop contextually relevant and culturally accurate datasets. Many existing African language datasets are derived from translations of high-resource languages, which can introduce misalignments and fail to capture authentic linguistic expressions. To improve NLP performance, datasets should be created using native speakers and sourced from real-world interactions, literature, and media in African languages. This will help preserve linguistic authenticity and avoid the overreliance on loanwords or structures that do not naturally occur in the language. Specialized corpora should also be developed to capture dialectal variations and culturally specific language use.
8 Conclusion
In this work, we set out to unravel how current African language policies, by dictating data availability and accessibility, directly influence model performance across the continent’s diverse linguistic spectrum. To achieve this, we conducted a comprehensive empirical evaluation of leading large language models (LLMs) on Sahara, an extensive benchmark assembled primarily from existing datasets. Sahara consists of $30$ publicly accessible datasets, covering $16$ classification, generation, MCCR, and token level tasks. We also introduced the first publicly available leaderboard for African languages. Our evaluations revealed critical gaps in task performance and language coverage, underscoring pronounced disparities rooted in policy-induced data inequities. Moreover, our findings highlight the importance of integrating theoretical insights on language policies with empirical evaluations to inform forward-thinking policy reforms and promote inclusive data practices. We believe that Sahara will serve as essential benchmarking standards, catalyzing future research and development efforts aimed at bridging the digital divide and fostering linguistic diversity in artificial intelligence for African communities.
Acknowledgments
We acknowledge support from Canada Research Chairs (CRC), CLEAR Global, the Natural Sciences and Engineering Research Council of Canada (NSERC; RGPIN-2018-04267), the Social Sciences and Humanities Research Council of Canada (SSHRC; 895-2020-1004; 895-2021-1008), Canadian Foundation for Innovation (CFI; 37771), Digital Research Alliance of Canada, [^2] https://alliancecan.ca and UBC ARC-Sockeye. [^3] https://arc.ubc.ca/ubc-arc-sockeye We also thank Clémentine Fourrier and Nouamane Tazi from HuggingFace for providing feedback on this work. The findings and conclusions contained within this work those of the authors and do not necessarily reflect positions or policies of any supporters.
9 Limitations
We identify the following limitations for our work:
  1. 1.
    Data and Data Availability Bias. Our benchmark relies on publicly available datasets, which inherently reflect existing biases in language documentation and institutional support. Many Indigenous African languages remain severely underrepresented due to historical policies that prioritize dominant languages. As a result, our findings may not fully capture the linguistic diversity of the continent.
  2. 2.
    Task Scope. The evaluation primarily focuses on tasks where data is available, such as language identification and text classification. However, the scarcity of datasets for complex NLP tasks (e.g., machine translation, question answering) limits our ability to assess model performance in broader applications. Future efforts should address this gap by developing more comprehensive datasets across a wider range of tasks.
  3. 3.
    Model Performance Constraints. While we analyze the impact of data scarcity on model effectiveness, our study does not explore architectural modifications or fine-tuning strategies that could improve performance for low-resource languages. Further research is needed to examine whether adaptive techniques, such as multilingual pretraining with targeted data augmentation, could mitigate these challenges.
  4. 4.
    Policy Implementation Challenges. Our recommendations for policy reforms and inclusive data practices require significant institutional commitment and resources. While we outline actionable steps, their adoption depends on the willingness of governments, research institutions, and technology developers to prioritize African language inclusion in NLP.
Despite these limitations, our work underscores the urgency of addressing data inequities and advocating for policies that foster greater linguistic diversity in AI development.
10 Ethical Considerations
We make the following ethics-related statements about our work:
  1. 1.
    Our research aims to investigate the progress of NLP for African languages, addressing historical language policies that have contributed to data inequities. By improving representation in NLP, we align with broader efforts to foster linguistic diversity in AI, ensuring that African languages are not sidelined in technological advancements.
  2. 2.
    The datasets used in our benchmark are derived from publicly available sources, yet their existence is shaped by historical and contemporary policies that prioritize certain languages over others. We acknowledge that the digital presence of African languages is not merely a technical issue but a policy-driven reality that influences which languages receive institutional support and investment.
  3. 3.
    Although we do not develop models in this work, our findings highlight the impact of policy-induced data disparities on model performance. Addressing these challenges requires policy interventions that support multilingual data collection, equitable resource allocation, and ethical considerations in dataset creation.
  4. 4.
    Proper attribution to dataset authors is not just an academic necessity but a policy imperative for transparency and recognition. To mitigate concerns about data ownership and ethical usage, we provide a publicly accessible reference file citing all datasets included in our benchmark. We strongly encourage researchers and policymakers to acknowledge these sources, reinforcing the importance of ethical and inclusive data practices.
Appendix A Sahara Task Clusters
A.1 Text Classification
Cross-Lingual Natural Language Inference.
This task involves the classification of 2 given sentences—a premise and a hypothesis into entailment, neutral, or contradiction semantic relation Adelani et al. (2024b). Adelani et al. (2024b) introduced AfriXNLI, a benchmark for evaluating 15 African Languages on Natural Language Inference (NLI). The languages were translated from the English portion of XNLI Conneau et al. (2018).
Language Identification.
We include language identification datasets across $518$ African languages Adebara et al. (2022). The dataset includes data from multiple domains including education, government, health care, news, and religion. Additional statistics of the data is available in Table D3. We include language identification tasks to investigate (1.) the capabilities of language models in correctly generating the appropriate language names of an input text and (2.) Correctly generate an additional sentence in a specific language. We include these two approaches because many languages have multiple names Chen et al. (2024) and a LM’s inability to appropriately name a language does not implicate a lack of information about the language.
News Classification.
This task involves automatically categorizing news articles into predefined categories or topics based on their content. The goal is to leverage machine learning and NLP techniques to understand and organize large volumes of news data, making it easier for users to access relevant information. News classification plays a crucial role in information retrieval, content recommendation, and other applications that involve organizing and categorizing textual data. The news classification cluster consists of data from four languages including Amharic Azime and Mohammed (2021), Kinyarwanda Niyongabo et al. (2020), Kirundi Niyongabo et al. (2020), and Swahili David (2020a, b). Table D3 provides additional details of the data in this cluster.
Sentiment Analysis.
Sentiment analysis is crucial in gaining insights into public opinion, customer feedback, and user sentiments across various platforms. The sentiment analysis cluster consists of the Bambara Sentiment dataset Diallo et al. (2021), YOSM–a Sentiment Corpus for Movie Reviews in Yorùbá Shode et al. (2022b), and the Nigerian Pidgin sentiment dataset Oyewusi et al. (2020), respectively. Further details about the statistical composition of this data is available in Table D3.
Topic Classification.
The primary goal for this task is to automatically categorize the content of the text based on its theme. Topic classification is widely used in various applications, such as document organization, content recommendation, and information retrieval. We include topic classification datasets for Yorùbá and Hausa Hedderich et al. (2020).
A.2 Text Generation
Machine Translation.
A robust benchmark for MT tasks is essential. Sahara, in its quest to effectively assess machine translation for African languages, selectively incorporates datasets containing diverse African languages into its benchmark. Specifically, Sahara includes datasets such as Pidgin-UNMT [^4] PidginUNMT GitHub Link Ogueji and Ahia (2019), Afro-MT [^5] AfroMT GitHub Link Reid et al. (2021a), Lafand-MT Adelani et al. (2022a), SALT [^6] SALT GitHub Link Akera et al. (2022), and HornMT, thus encompassing a total of $45$ language pairs.
Paraphrase.
Paraphrasing is a core task in NLG that revolves around the comprehension and generation of text that conveys the same meaning while possibly exhibiting differences in structure, phrasing, or style Chen et al. (2023). Sahara incorporates the TaPaCo dataset Scherrer (2020), which is publicly accessible and covers a total of $73$ languages, including four of African origin. These African languages consist of Kirundi, Amazigh, a macro-language Berber, and Afrikaans.
Summarization.
Summarization is the task of creating a condensed version of a text that retains its core ideas and essential information Nallapati et al. (2016). In the context of Sahara, our focus primarily centers on a subset of the abstractive summarization dataset, XL-SUM Hasan et al. (2021), which includes several African languages: Amharic, Hausa, Igbo, Kirundi, Oromo, Pidgin, Somali, Swahili, Tigrinya, and Yoruba. Additionally, we have extended our dataset collection efforts by scraping non-overlapping summarization datasets from BBC and Voice of Africa (VoE) websites, specifically targeting three African languages: Hausa, Ndebele, and Swahili.
Title Generation.
The task of title generation in NLG involves generating a concise and meaningful title or headline for a given article. This task shares its dataset source with summarization, drawing from XL-SUM, and also includes dataset extensions from BBC and VoE with focus on enabling zero-shot title generation.
A.3 Multiple-Choice, Comprehensive and Reasoning
This task cluster majorly addresses Question Answering type questions. The QA tasks in this cluster are more mathematical-reasoning and reading-comprehension-oriented.
Mathematical Word Problems.
Adelani et al. (2024b) introduced AfriMGSM, a QA dataset that contains human-written grade school mathematical word problems. Their dataset covers 15 African Languages, and the translation of each question to English for each language for in-language evaluation of LLMs. In this task, given a word problem, the LLM is expected to return a numerical value as output.
General Knowledge QA.
Adelani et al. (2024b) introduced AfriMMLU, a multi-choice QA dataset covering 5 areas of knowledge in 15 African Languages. The knowledge areas are high-school geography, high-school microeconomics, elementary mathematics, international law), and global facts. In this task given a question in any of the aforementioned areas, and 4 options, the LLM is expected to choose the right option.
Reading Comprehension QA.
In this task, the LLMs are evaluated based on their ability to understand information in a given article. Bandarkar et al. (2024b) introduced a parallel reading comprehension dataset covering 122 languages, 23 of which are native to Africa. The dataset comprises of short paragraphs accompanied by multi-choice Questions, whose answers can be inferred from accompanying paragraphs.
Context-based Question-Answering.
QA represents another vital facet of NLG, focusing on a model’s capability to provide answers based on integrated knowledge. Sahara relies on the gold passage dataset from the multilingual TYDIA [^7] TyDiQA GitHub Link QA resource Clark et al. (2020). This dataset offers concise questions with precisely one corresponding answer, enabling rigorous evaluation of QA performance.
A.4 Tokens-level Classification
Named Entity Recognition.
We incorporate NER datasets across $30$ languages. We use Distance Supervision NER (DS NER) Data Hedderich et al. (2020), MasakhaNER Adelani et al. (2022b), WikiAnn Rahimi et al. (2019), Yoruba-Twi NER Alabi et al. (2020b), and multiple NER data from SADiLaR. Additional details about the datasets are available in Table D3.
Phrase Chunking.
This task involves identifying and grouping together consecutive words or tokens into meaningful syntactic units, known as phrases. These phrases can include noun phrases, verb phrases, prepositional phrases, and other grammatical structures. The goal of phrase chunking is to analyze the grammatical structure of a sentence and extract higher-level syntactic units for better understanding. The phrase chunking cluster consists of data for ten Indigenous languages of South Africa (see Table D3). The data has annotations for noun, verb, adjective, adverbial, and prepositional phrase chunks. Words not belonging to these phrase types are labelled with the tag O.
Part of Speech Tagging.
This classification task involves assigning grammatical categories or labels (such as noun, verb, adjective) to each word in a sentence. The primary goal of POS tagging is to analyze the syntactic structure of a sentence and understand the role of each word within it. POS tagging is crucial for various NLP applications, as it provides a foundation for more advanced linguistic analysis. We include datasets for Igbo taken from IgboNLP Onyenwe et al. (2019).
Appendix B Benchmark
Table B1 shows the comparison between Sahara with others
Category Benchmark Reference Lang/Total Tasks
Multilingual FLoRES200 Costa-jussà et al. (2022) 52/200 1
GEMv1 Gehrmann et al. (2021) 10/52 5
GEMv2 Gehrmann et al. (2021) 10/52 9
NLLB M.D. Costa-jussà et al. (2022) 2/8 1
NLLB S.D. Costa-jussà et al. (2022) 2/8 1
SIB-200 Adelani et al. (2024a) 46/200 1
Toxicity200 Costa-jussà et al. (2022) 50/200 1
XGLUE Liang et al. (2020) 1/19 3
XTremE Hu et al. (2020) 2/40 4
Language-specific BanglaNLPv1 Bhattacharjee et al. (2023) 1 6
CLUE Xu et al. (2020) 1 6
CUGE Yao et al. (2021) 1 8
Dolphin Nagoudi et al. (2023) 1 13
FLUE Le et al. (2020) 1 5
IndoNLU Wilie et al. (2020) 1 5
IndoNLG Guntara et al. (2020) 1 6
JGLUE Kurihara et al. (2022) 1 3
KorNLU Ham et al. (2020) 1 2
LOT Guan et al. (2022) 1 4
ORCA Elmadany et al. (2023) 1 7
PhoMT Doan et al. (2021) 1 1
African AfriSenti Muhammad et al. (2023a) 14 1
AfroMT Reid et al. (2021b) 8/8 1
AfroNLG Adebara et al. (2024b) 517/517 6
AfroNLU Adebara et al. (2023) 28/28 7
Horn-MT -- 3/3 1
Iroko-Bench Adelani et al. (2024b) 16/16 3
Mafand-MT Adelani et al. (2022a) 17/17 1
Menyo-20k Adelani et al. (2021b) 1/1 1
Naija-senti Muhammad et al. (2022) 4/4 1
[Uncaptioned image] Sahara ours 517/517 16
Table B1: A Comparison of Sahara with other Benchmarks. M.D: Multilingual domain, S.D: Seed Data.
Appendix C Few-shot prompt example
Figure C1 shows an example for MT task.
Figure C1: Few-shot prompting example for the machine translation task. The prompt provides five examples of the specific task for context and alignment.
Appendix D Dataset Description
Table D2 describes all datasets included in Sahara
Dataset Brief Description African Language(s)
MasakhaNER Adelani et al. (2021a), Adelani et al. (2022c) Curated for Named Entity Recognition for several African Languages amh, bam, bbj, ewe, fon, hau, ibo, kin, lug, luo, mos, nya, pcm, sna, swa, tsn, twi, wol, xho, yor, zul
MasakhaPOS Dione et al. (2023a) Curated from News sources and publicly available datasets for Part-Of-Speech Tagging bam, bbj, ewe, fon, hau, ibo, kin, lug, luo, mos, nya, pcm, sna, swa, tsn, twi, wol, xho, yor, zul
MAFAND-MT Adelani et al. (2022a) African news corpus for Machine Translation bam, bbj, ewe, fon, hau, ibo, lug, luo, mos, swa, tsn, twi, wol, xho, yor, zul
XL-Sum Hasan et al. (2021) Curated for Summarization and Title Generation amh, ibo, orm, run, swa, yor, hau, pcm, som, tir
YorubaSenti Orimaye et al. (2012) Curated from comments on movies on YouTube yor
TaPaCo Scherrer (2020) Developed to provide example sentences and translations for particular linguistic constructions and words afr, ber, run , zgh
TyDi QA Clark et al. (2020) Curated for Information-Seeking QA tasks swa
HornMT Parallel MT dataset curated for languages in Horn of Africa aaf, amh, orm, som, tir
yoruba-twi-ner Alabi et al. (2020a) Curated for evaluating embeddings for low-resource languages yor, twi
KINNEWS and KIRNEWS Niyongabo et al. (2020) Curated for multi-class classification of news articles kin, run
AmharicNews Azime and Mohammed (2021) Curated for text classification from news articles amh
SwahiliNews v0.2 David (2020a) Curated for text classification from news articles swa
Bambara Diallo et al. (2021) Curated for Sentiment Analysis through web crawling bam
IgboNLPOnyenwe et al. (2018b) Contains text corpus, POS tagset and POS-tagged subcorpus ibo
yosm Shode et al. (2022a) Curated from movie reviews for Sentiment Analysis yor
pidgin-tweet Oyewusi et al. (2020) Curated for SA from tweets pcm
WikiAnn Pan et al. (2017), Rahimi et al. (2019) afr, amh, ibo, mlg, kin, som, swa, yor
HausaTopic,YorubaTobic Hedderich et al. (2020) Curated for Topic Classification hau, yor
SaDiLAR Eiselen (2016) Curated for NER of 10 South African Languages afr, nbl, nso, sot, ssw, tsn, tso, ven, xho, zul
AfriSenti Muhammad et al. (2023a) Curated for African tweet sentiment analysis amh, ALG-ara, hau, kin, ibo, MOR-ara, MOZ-por, pcm,form, swa, tir, twi,tso,yor
Afro-MT Reid et al. (2021a) Benchmark for MT of 8 African Languages afr, bem, lin, run, sot, swa, xho, zul
PidginUNMT Ogueji et al. (2021) pcm
SALT Akera et al. (2022) Curated for Parallel MT of 5 Ugandan Languages ach, lgg, lug, nyn, teo
Cheetah Adebara et al. (2024b) Summarization test set hau, nde, swa
Cheetah Adebara et al. (2024b) Title generation test set amh, gag (zero-shot), hau, ibo, pcm, som, swa, tir, yor, kin (zero-shot), afr, mlg (zero-shot), orm, nde (zero-shot), sna(zero-shot)
NCHLT-NER Eiselen (2016) Named Entity Recognition afr, nbl, nso, sot, ssw, tsn, tso, ven, xho, zul
BeleBele Bandarkar et al. (2024b) Reading Comprehension MCQA afr amh, bam, fuv, gaz, hau, ibo,kea,kin, lin, lug, luo,nso, nya, sna ,som, sot, ssw, swh, tir, tsn, tso, wol, xho, yor,zul
IrokoBench Adelani et al. (2024b) MGSM,MMLU,XNLI amh, ewe, hau, ibo,kin, lin, lug, orm, sna ,sot, swa, tir, wol, xho, yor,zul
Table D2: A description of all datasets included in Sahara. ALG: Algerian, MOR: Moroccan, MOZ: Mozambican.
Task Source Language Train/ Dev/ Test
MT Adelani et al. (2022a) eng-hau 5.9K/ 1.3K/ 1.5K
eng-ibo 6.9K/ 1.5K/ 1.4K
eng-lug 4.1K/ 1.5K/ 1.5K
eng-pcm 4.8K/ 1.5K/ 1.6K
eng-swa 30.8K/ 1.8K/ 1.8K
eng-tsn 2.1K/ 1.3K/ 1.5K
eng-twi 3.3K/ 1.3K/ 1.5K
eng-yor 6.6K/ 1.5K/ 1.6K
eng-zul 3.5K/ 1.5K/ 1.0K
fra-bam 3.0K/ 1.5K/ 1.5K
fra-bbj 2.2K/ 1.1K/ 1.4K
fra-ewe 2.0K/ 1.4K/ 1.6K
fra-fon 2.6K/ 1.2K/ 1.6K
fra-mos 2.5K/ 1.5K/ 1.6K
fra-wol 3.4K/ 1.5K/ 1.5K
Reid et al. (2021a) eng-afr 25.8K/ 3.2K/ 3.2K
eng-bem 12.0K/ 1.5K/ 1.5K
eng-lin 17.7K/ 2.2K/ 2.2K
eng-run 12.5K/ 1.6K/ 1.6K
eng-sot 28.8K/ 3.6K/ 3.6K
eng-swa 28.1K/ 3.5K/ 3.5K
eng-xho 26.1K/ 3.3K/ 3.3K
eng-zul 29.1K/ 3.6K/ 3.6K
Ogueji and Ahia (2019) eng-pcm 1.7K/ 0.2K/ 0.2K
Akera et al. (2022) All-pairs1 20.0K/ 2.5K/ 2.5K
HornMT All-pairs2 1.7K/ 0.2K/ 0.2K
NER Adelani et al. (2021a) See Table D2 41.2K/ 5.1K/ 5.1K
Adelani et al. (2022c) See Table D2 41.2K/ 5.1K/ 5.1K
Eiselen (2016) See Table D2 1.7M/ 219.7K/ 215.6K
Alabi et al. (2020a) See Table D2 20.2K/ 2.8K/ 5.5K
Pan et al. (2017) See Table D2 9.2K/ 9.2K/ 9.4K
News Azime and Mohammed (2021) amh 41.2K/ 5.1K/ 5.1K
Niyongabo et al. (2020) kin 15.3K/ 1.7K/ 4.3K
run 3.3K/ 0.4K/ 0.9K
David (2020a) swa 4.4K/ 0.6K/ 0.6K
Paraphrase Scherrer (2020) amh, ber, kab, run 22.4K/ 2.8K/ 2.8K
ber 17.6K/ 2.2K/ 2.2K
kab 4.4K/ 0.6K/ 0.6K
Phrase Chunking Eiselen (2016) See Table D2 107.5K/ 13.0K/ 13.4K
POS Tagging Onyenwe et al. (2018a, 2019) ibo 756.8K/ 94.7K/ 95.0K
Question Answering Clark et al. (2020) swa 49.9K/ 0.5K/ n/a
Sentiment Analysis Diallo et al. (2021) bam 2.4K/ 0.3K/0.3K
Oyewusi et al. (2020) pcm 11.2K/ 1.4K/ 1.4K
Shode et al. (2022a) yor 0.8K/ 0.2K/ 0.5K
Summarization Hasan et al. (2021) amh 5.8K/ 0.7K/ 0.7K
ibo 4.2K/ 0.5K/ 0.5K
orm 6.1K/ 0.8K/ 0.8K
Rundi 5.7K/ 0.7K/ 0.7K
swa 7.9K/ 1.0K/ 1.0K
yor 6.4K/ 0.8K/ 0.8K
hau 6.4K/ 0.8K/ 0.8K
pcm 9.2K/ 1.2K/ 1.2K
som 6.0K/ 0.7K/ 0.7K
Tigrinya 5.5K/ 0.7K/ 0.7K
[…]
Title Generation Hasan et al. (2021) amh 5.8K/ 0.7K/ 0.7K
ibo 4.2K/ 0.5K/ 0.5K
orm 6.1K/ 0.8K/ 0.8K
run 5.7K/ 0.7K/ 0.7K
swa 7.9K/ 1.0K/ 1.0K
yor 6.4K/ 0.8K/ 0.8K
hau 6.4K/ 0.8K/ 0.8K
pcm 9.2K/ 1.2K/ 1.2K
som 6.0K/ 0.7K/ 0.7K
Tigrinya 5.5K/ 0.7K/ 0.7K
Adebara et al. (2024b) $\star$ -/ -/ 5.9K
Topic Classification Hedderich et al. (2020) hau 2.0K/ 0.3K/0.6K
yor 1.3K/ 0.2K/ 0.4K
[…]
Appendix E Data-Performance-Policy Analysis
(a) Languages with an F1 score of $50$ or higher.
(b) Languages with an F1 score below $50$.
Figure E2: Overall distribution of Claude-3.5-Sonnet’s performance on language identification.
Figure E2 shows the overall distribution of Claude-3.5-Sonnet’s (best model) performance on language identification. In sub-figure 2(a), the vertical axis represents the F1 score, and each point along the horizontal axis corresponds to a specific African language for which the model achieves an F1 score of $50$ or higher. Only around $50$ languages fall into this category, with their performance ranging from just above $50$ F1 to approximately $90$ F1. Meanwhile, sub-figure 2(b) depicts the majority of African languages, all of which register F1 scores below $50$. These results indicate that while the model manages moderate or better performance on a limited subset of languages, it struggles substantially with the remaining ones. Overall, this visualization underscores that the model’s ability to correctly identify and process most African languages remains limited, highlighting a need for more comprehensive data coverage and targeted model improvements.